The lateral area of a right cylindrical tank is 48πm2

 and its volume is 72πm3.

 Find the altitude of the tank.


1
Expert's answer
2022-02-28T14:26:33-0500

The lateral area is given by "S=2\pi r h" while his volume is given by "V=\pi r^2 h". We remark that the ratio "\frac{S^2}{V}= \frac{4\pi^2 r^2 h^2}{\pi r^2 h}=4\pi h". This gives us "h=\frac{S^2}{4\pi V}=\frac{48^2 \pi^2}{4\pi^2\cdot 72}=8m"